OpenStudy (anonymous):

The mass defect of a particular reaction resulted in the production of 5.50 x 1015 joules of energy. What mass (if totally converted to energy) would correspond to this amount of energy? (1 J = 1 kg m2/s2) (Points : 3) 1.83 x 108 kg 4.95 x 107 kg 6.11 x 10-2 kg 1.64 x 10-1 kg

OpenStudy (dean.shyy):

Take a look at this: http://is.gd/a1iAPZ

OpenStudy (anonymous):

I got 1.83 x 10^8 kg-- the first one

OpenStudy (anonymous):

here again,use einstein's mass-energy equation.your answer should be 6.11*10^-2 kg

OpenStudy (anonymous):

E=mc^2 (5.5x10^15 kg m^2/s^2)=mass*(3x10^8 m/s)^2 solve for mass